A chemist needs 160 milliliters of a 36% solution but has only 24% and 72% solutions available. Find how many milliliters of each that should be mixed to get the desired solution.
A) 40 ml of 24%; 120 ml of 72%
B) 123 ml of 24%; 40 ml of 72%
C) 120 ml of 24%; 40 ml of 72%
D) 123 ml of 24%; 37 ml of 72%

Answer by ankor@dixie-net.com(22740)

A chemist needs 160 milliliters of a 36% solution but has only 24% and 72%
solutions available.
Find how many milliliters of each that should be mixed to get the desired solution.
:
Let x = amt of 72% solution required
The resulting amt is to be 16 ml, therefore
(160-x) = amt of 24% solution
:
A typical mixture equation using the decimal equiv of per cent
:
.72x + .24(160-x) = .36(160)
.72x + 38.4 - .24x = 57.6
.72x - .24x = 57.6 - 38.4
.48x = 19.2
x = 19.2/.48
x = 40 ml of 72% solution required
:
I'll let you find the amt of 24% required